LINE INTENSITIES OF PHOSPHINE IN THE 10 μm REGION FOR PLANETARY SPECTRA ANALYSIS.

L. R. BROWN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109, USA; R. L. SAMS, Pacific Northwest National Laboratory, P.O. Box 999, Mail Stop K8-88, Richland, WA 99352, USA; I. KLEINER, C. COTTAZ, Laboratoire de Photophysique Moléculaire, Université Paris Sud, 91405 Orsay Cedex, France.

A preliminary study of the line intensities of phosphine PH_{3} in the 10μm region between 800 and 1300 cm^{−1} is reported. The objective is to provide a prediction for planetary spectra studies of the absorption in the Saturnian atmosphere. Over 1000 intensity measurements were performed using the Fourier transform spectrometer located at the Pacific Northwest Laboratory at a resolution of 0.0011 cm^{−1}. Using the rovibrational line assignments and positions published by Fusina and Di Lonardo^{a} for the two fundamental bands ν_{2} and ν_{4} located near 992.13 and 1118.31 cm^{−1} respectively, and treating these two bands within a dyad interacting system, we will report preliminary intensity analysis for this spectral range.^{b}

aL. Fusina and G. Di Lonardo J. Mol. Struct. \underline{517-518}, 67 (2000).

bPart of the research reported in this paper was performed at the Jet Propulsion Laboratory, California Institute of Technology, under contract with the National Aeronautics and Space Administration.